# Crossing the Human-Robot Embodiment Gap with Sim-to-Real RL using One Human Demonstration

### v1

@@ I Introduction
Figure 1: Human2Sim2Robot is a real-to-sim-to-real RL framework for learning dexterous manipulation policies from one human RGB-D video demonstration per task. From the demonstration, we leverage object pose trajectories and pre-manipulation hand poses to provide guidance and advantageous initialization for RL policy training in simulation. These policies demonstrate zero-shot sim-to-real transfer to a Kuka arm and Allegro hand.
@@ III-A Real-to-Sim & Human Demonstration
Lastly, we retarget the human pre-manipulation hand pose to a robot hand pose through a two-step IK procedure using cuRobo [48]. Figure 3 visualizes how we perform human-to-robot hand retargeting. In the first step, the robot arm’s joint angles are adjusted to align the base position and orientation of the robot hand’s middle knuckle with that of the human hand (with a small offset, see Appendix -B for details). In the second step, the robot hand’s joint angles are adjusted to align the robot’s fingertip positions with the corresponding human fingertip positions. This generates a pre-manipulation robot hand pose $(\bm{T}^{\text{wrist}},\bm{q}^{\text{hand}})$ for the object pose represented as $\bm{T}^{\text{goal}}_{\tau}$, where $\bm{T}^{\text{wrist}}\in\text{SE}(3)$ is the robot wrist pose, and $\bm{q}^{\text{hand}}\in\mathbb{R}^{N^{\text{hand-joints}}}$ is robot hand joint configuration. $N^{\text{hand-joints}}=23$ for the Allegro hand. This simple IK process faithfully retargets the human pre-manipulation hand pose, while maintaining kinematic feasibility and alignment between human and robot hand configurations.
@@ III-B Simulation-based Policy Learning
$\centering\bm{o}_{t}=[\bm{q}_{t},\bm{\dot{q}}_{t},\bm{X}^{\text{fingertips}}_{t},\bm{x}^{\text{palm}}_{t},\bm{X}^{\text{obj}}_{t},\bm{X}^{\text{goal}}_{t}]\@add@centering$ (3)
where $\bm{q}_{t}\in\mathbb{R}^{N^{\text{joints}}}$ is the vector of robot joint angles, $\bm{\dot{q}}_{t}\in\mathbb{R}^{N^{\text{joints}}}$ is the vector of robot joint velocities, $\bm{X}^{\text{fingertips}}_{t}\in\mathbb{R}^{N^{\text{fingers}}\times 3}$ contains 3D positions of each fingertip, $\bm{x}^{\text{palm}}_{t}\in\mathbb{R}^{3}$ is the palm center’s 3D position, and $\bm{X}^{\text{obj}}_{t},\bm{X}^{\text{goal}}_{t}\in\mathbb{R}^{N^{\text{anchor}}\times 3}$ contain 3D positions of the anchor points representing the current and goal object pose, respectively. $N^{\text{joints}},N^{\text{fingers}}\in\mathbb{N}$ are the number of robot joints and fingers, respectively.
@@
We use an initial state distribution, guided by the human pre-manipulation hand pose, to simplify exploration and bias the policy toward learning human-like behavior. We construct this distribution by sampling the object pose around the trajectory’s initial object pose, then set the robot configuration to match the pre-manipulation hand pose with slight perturbation (see Appendix -E for details). Given this perturbed hand pose, we compute a feasible arm joint angle configuration with IK  [^1] We use cuRobo [48] to perform parallelized, collision-free IK to ensure RL training performance is not bottlenecked by IK computations. The environment is reset if the object is too far from the goal $d(\bm{T}^{\text{goal}}_{t},\bm{T}^{\text{obj}}_{t})>D_{\text{max}}$, the robot palm is too far from the object $d(\bm{T}^{\text{palm}}_{t},\bm{T}^{\text{obj}}_{t})>D_{\text{max}}$, or the total timesteps of the goal trajectory is reached $\tau+t>T$, where $D_{\text{max}}=0.25m$.

### v2

@@ I Introduction
Figure 1: Our Framework. Human2Sim2Robot is a real-to-sim-to-real RL framework for learning dexterous manipulation policies from one human RGB-D video demonstration per task. From the demonstration, we leverage object pose trajectories and pre-manipulation hand poses to provide guidance and advantageous initialization for RL training in simulation. These policies demonstrate zero-shot sim-to-real transfer to a Kuka arm and Allegro hand.
@@ III-A Real-to-Sim & Human Demonstration
Lastly, we retarget the human pre-manipulation hand pose to a robot hand pose through a two-step IK procedure using cuRobo [48]. Figure 3 visualizes how we perform human-to-robot hand retargeting. In the first step, the robot arm’s joint angles are adjusted to align the base position and orientation of the robot hand’s middle knuckle with that of the human hand (with a small offset, see Appendix -B for details). In the second step, the robot hand’s joint angles are adjusted to align the robot’s fingertip positions with the corresponding human fingertip positions. This generates a pre-manipulation robot hand pose $(\bm{T}^{\text{wrist}},\bm{q}^{\text{hand}})$ for the object pose represented as $\bm{T}^{\text{goal}}_{\tau}$, where $\bm{T}^{\text{wrist}}\in\text{SE}(3)$ is the robot wrist pose, and $\bm{q}^{\text{hand}}\in\mathbb{R}^{N^{\text{hand-joints}}}$ is robot hand joint configuration. $N^{\text{hand-joints}}=16$ for the Allegro hand. This simple IK process faithfully retargets the human pre-manipulation hand pose, while maintaining kinematic feasibility and alignment between human and robot hand configurations.
@@ III-B Simulation-based Policy Learning
$\centering\bm{o}_{t}=[\bm{q}_{t},\bm{\dot{q}}_{t},\bm{X}^{\text{fingertips}}_{t},\bm{x}^{\text{palm}}_{t},\bm{X}^{\text{obj}}_{t},\bm{X}^{\text{goal}}_{\tau+t}]\@add@centering$ (3)
where $\bm{q}_{t}\in\mathbb{R}^{N^{\text{joints}}}$ is the vector of robot joint angles, $\bm{\dot{q}}_{t}\in\mathbb{R}^{N^{\text{joints}}}$ is the vector of robot joint velocities, $\bm{X}^{\text{fingertips}}_{t}\in\mathbb{R}^{N^{\text{fingers}}\times 3}$ contains 3D positions of each fingertip, $\bm{x}^{\text{palm}}_{t}\in\mathbb{R}^{3}$ is the palm center’s 3D position, and $\bm{X}^{\text{obj}}_{t},\bm{X}^{\text{goal}}_{\tau+t}\in\mathbb{R}^{N^{\text{anchor}}\times 3}$ contain 3D positions of the anchor points representing the current and goal object pose, respectively. $N^{\text{joints}},N^{\text{fingers}}\in\mathbb{N}$ are the number of robot joints and fingers, respectively.
@@
We use an initial state distribution, guided by the human pre-manipulation hand pose, to simplify exploration and bias the policy toward learning human-like behavior. We construct this distribution by sampling the object pose around the trajectory’s initial object pose, then set the robot configuration to match the pre-manipulation hand pose with slight perturbation (see Appendix -E for details). Given this perturbed hand pose, we compute a feasible arm joint angle configuration with IK  [^1] We use cuRobo [48] to perform parallelized, collision-free IK to ensure RL training performance is not bottlenecked by IK computations. The environment is reset if the object is too far from the goal $d(\bm{T}^{\text{goal}}_{\tau+t},\bm{T}^{\text{obj}}_{t})>D_{\text{max}}$, the robot palm is too far from the object $||\bm{x}^{\text{palm}}_{t}-\bm{x}^{\text{obj}}_{t}||>D_{\text{max}}$, or the total timesteps of the goal trajectory is reached $\tau+t>T$, where $D_{\text{max}}=0.25m$.